Projection Admin 3/17

To do this evening/ was completed: 

Presentation Issues:

Preshow Clips/Playlist: KGO playlist built, Watership Down playlist built/tech'd and REC playlist updated with credits/volume info. Subs look great. - L

Adjusted and scheduled the Full Metal Jacket pre-show playlist in theater 1 -WB

Trailer Updates:

Special Events/ AV set up: REMINDER for tomorrow:  Superhuman Body screening (tomorrow, in T1, @10am again!) built the playlist and scheduled it on the TMS, so it's all good to go.  Checked to see if the babies playlist for OCAP Mickey 17 had been scheduled and it had not, so made that one and scheduled it. Please remember the babies screenings! 
Also, playlist for the school screening on the 27th is built and scheduled. Snow White.- L

Ordering/Receiving (posters & trailers): Received and stored posters for Bring Her Home -WB

Ordering/Receiving (content): Reached out to Deluxe for the keys to Tyrone, which will be coming in later tonight. 
The DCP of Paper Moon gets dropped off tomorrow. 
Downloaded, DCP'd and ingested Kevin's special preshow for KGO. 
Reached out for an eta for Zola and we should be receiving it by satellite tomorrow afternoon.  
Key delivery for Snow White will be tonight,  Alto Knights tomorrow! - L

Outgoing/Shipping (content): 

Print/video Inspection: Ran house checks in all auditoriums, tech'd the Gremlins 2 pre-show and adjusted and added volume cues across all first run and special event SPLs where necessary (the Kevin Geeks Out and Conquest trailers were a tad loud at 4.0 + added an 8.0 cue for the start of the 35mm) -WB

Got eyes on the intro slide for They Cloned Tyrone (keys for the feature will not be made available until tonight so this content can be tech'd tomorrow during admin hours) and it looks great -WB

Asset tracking doc updated/checked?:

Maintenance/cleaning: T2 is the only lamp we'll need to replace somewhat soon - a little under 2 weeks until the warning. Storage space is solid. Prepped update and info for Wednesday, trying to get ahead of the calendar with upcoming events/coverage needed. - L

Swept and mopped booth 1 ahead of tonight's 35mm screening -WB

Cue sheets updated/checked (every shift): Yes. Sheet checked, Gremlins checked. 

Marquee updated?:

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Open 12/28

Projection Admin (and Close) 10/19

NPP Projection Report Friday 1/31 CLOSE Chris